Onshore Bank has $28 million in assets, with risk-adjusted assets of $18 million. Core Equity Tier 1 (CET1) capital is $950,000, additional Tier I capital is $210,000, and Tier II capital is $416,000. The current value of the CET1 ratio is 5.28 percent, the Tier I ratio is 6.44 percent, and the total capital ratio is 8.76 percent.

A. Calculate the new value of CET1, Tier I, and Total capital ratios for the following transactions: The bank issues $2.8 million of CDs and uses the proceeds to issue category 1 mortgage loans with a loan-to-value ratio of 80 percent.

B. Calculate the new value of CET1, Tier I, and Total capital ratios for the following transactions: The bank repurchases $108,000 of common stock with cash .

C.Calculate the new value of CET1, Tier I, and Total capital ratios for the following transactions: The bank receives $508,000 in deposits and invests them in T-bills.
Answer #1

A. Issue of CDs will not affect CET1, Tier I and Total Capital. It will affect bank's liabilities. Issue of mortgage loans with a loan-to-value ratio of 80 percent will increase the risk-adjusted figure by $2.24 million ($2.8 million * 80%).

CET1 = $950,000/($18 million + $2.24 million) = 4.69%

Tier 1 ratio = ($950,000+$210,000)/($18 million + $2.24 million) = 5.73%

Total Capital Ratio = ($950,000+$210,000+$416,000)/($18 million + $2.24 million) = 7.79%

B. Bank repurchases $108,000 of common stock with cash. It will reduce our Core Equity Tier 1 with $108,000. Risk-adjusted asset will not reduce because risk weight of cash is 0.

CET1 = ($950,000-$108,000)/($18 million) = 4.68%

Tier 1 ratio = ($950,000-$108,000+$210,000)/($18 million) = 5.84%

Total Capital Ratio = ($950,000-$108,000+$210,000+$416,000)/($18 million) = 8.15%

C. Receipt of $508,000 in deposits will not affect capital and will be treated as liabilities. Investment of that amount in T-bills will also not affect risk-adjusted asset reason being risk weight is 0. So this transaction will not affect CET1,Tier I and Total capital ratio

CET1 = $950,000/($18 million) = 5.28%

Tier 1 ratio = ($950,000+$210,000)/($18 million) = 6.44%

Total Capital Ratio = ($950,000+$210,000+$416,000)/($18 million) = 8.76%